# Target Specific De Novo Design of Drug Candidate Molecules with Graph Transformer-based Generative Adversarial Networks

### v6

@@ 2.3 Ablation Study
In this analysis, we compare the output of different DrugGEN model variations with each other and with the molecules in the training datasets. We used the same set of fundamental benchmarking metrics as in the previous section, as well as the synthetic accessibility (SA) metric [43] (see “Performance metrics” for details). For this experiment, we utilised four different models: (i) DrugGEN: the default model, (ii) DrugGEN-NoTarget: a DrugGEN model that only generates random molecules, which has the same architecture as DrugGEN; however, its discriminator takes random real molecules as input instead of target specific inhibitors (the sole aim is to generate drug-like molecules), (iii) MLP baseline: a targeted baseline model with multi-layered perceptron (MLP) based generator/discriminator modules, and (iv) MLP-NoTarget baseline: a MLP-based non-targeted baseline model. These baseline models are differentiated from DrugGEN by replacing the graph transformer encoder blocks with simple MLP layers (please see “Methods” for details). The purpose behind comparing the main DrugGEN model and its non-targeted version with the MLP-based baselines is to underscore the benefits of employing graph transformers for molecule generation. Targeted and non-targeted baseline models follow the same training schedule and datasets as DrugGEN and DrugGEN-NoTarget, respectively. We generated 10,000 de novo molecules from each model and ran them through the metrics mentioned above.
Table 2 shows ablation performance results, in which all models demonstrate high validity, uniqueness, novelty, and IntDiv scores, indicating the high efficiency of our molecular design system. Low SA and high QED scores pinpoint the easy synthesizability of the generated molecules, in comparison to ChEMBL data (random drug-like molecules) and real AKT1 inhibitors. DrugGEN’s generator module takes random molecules as the initial input (i.e., starting molecules). Therefore, measuring a second novelty score against the starting molecules used during the generation process is also essential. We named this metric novelty at inference (NI) and provided the scores in Table 2 in the same column as the regular novelty measure (i.e., novelty against training data). We observed that NI scores are slightly lower than the standard novelty values, which was expected since the input of the inference run has a more direct effect on generated samples. It is also important to note that the NI scores of DrugGEN are still higher than the regular novelty scores of most of the models listed in Table 1. Meanwhile, both the DrugGEN-NoTarget and the MLP-NoTarget baseline models scored low NI values (0.552 and 0.201, respectively), indicating that these models generate their starting molecules at least half of the time. Additional comments can be found in Supplementary Material (S3).
We also docked molecules belonging to each model/dataset to the AKT1 protein structure (details are provided in Methods). The median scores (using the best 10% of molecules in terms of docking scores) are shown in the docking score column of Table 2. The results indicate that the real AKT1 inhibitors obtained the best performance (-9.406 kcal/mol), which was expected since these are experimentally verified binders and included as positive control. The DrugGEN model has obtained the closest results to real AKT1 inhibitors (-8.369 kcal/mol). On the other hand, the MLP baseline model falls short (-6.367 kcal/mol), pointing out that MLP lacks the required generalisation capability to understand long-term relationships between atoms of the molecule to design effective inhibitors. On the other hand, the transformer architecture leverages the attention mechanism to effectively extract the intrinsic structural patterns. A more detailed investigation of the docking analysis is provided below, in the “Molecular docking analysis” section.
Overall, the ablation results indicate a better targeted generation performance when the graph transformer blocks are utilised. Based on the results in Tables 1 and 2, it is possible to state that DrugGEN generates valid, diverse and novel molecules, which docks well to the intended target.
Table 2: Ablation study performance results in terms of (i) fundamental benchmarking metrics (i.e., validity, novelty, uniqueness, and IntDiv), (ii) metrics measuring physicochemical and drug-likeness-related properties (i.e., QED, and SA), and (iii) docking scores (kcal/mol) obtained against the AKT1 protein structure (i.e., median of the scores of the top 10% of molecules). Arrows show the direction in which the performance increases.
Models/ Datasets Highlight of the Model/Dataset Validity ($\uparrow$) Novelty (NI*) ($\uparrow$) Uniqueness ($\uparrow$) IntDiv ($\uparrow$) QED ($\uparrow$) SA ($\downarrow$) Docking score ($\downarrow$) ChEMBL Data Training data (random mol.*) - - - 0.877 0.543 3.002 -6.871 Real AKT1 Inhibitors Training data (targeted mol.) - - - 0.827 0.460 3.051 -9.406 DrugGEN Targeted gen. via graph trans. 0.931 0.991 (0.949) 1.000 0.875 0.520 2.842 -8.369 MLP Baseline Targeted gen. via MLPs 0.934 0.992 (0.952) 1.000 0.860 0.519 2.833 -6.367 DrugGEN NoTarget Non-targeted gen. via graph trans. 0.913 0.990 (0.552) 1.000 0.884 0.578 3.030 - MLP-NoTarget Baseline Non-targeted gen. via MLPs 0.957 0.999 (0.201) 0.998 0.883 0.576 2.821 - *mol: molecules, NI: Novelty at Inference (novelty measured against the real molecules used as input at inference), MLP: multi-layered perceptron.
2.4 Physicochemical Property Comparisons
We drew histograms (density plots) displaying the physicochemical property value distributions of real molecules in our training dataset and de novo generated molecules from DrugGEN models and other target-based generation models (Figure S2). Details of this analysis are given in Supplementary Material (S4). In summary, de novo molecules belonging to the DrugGEN model and real AKT1 inhibitors share a similar region of the physicochemical property space in general (i.e., DrugGEN’s distribution shape and mean more closely resemble those of real AKT1 inhibitors compared to other models). This finding indicates that DrugGEN has effectively learned the physicochemical properties of the real inhibitors of the selected target protein. It is also important to note that this correlation is not due to merely overfitting to the training dataset, indicated by the high novelty scores (Table 1 and 2).
2.5 Molecular Docking of de novo Molecules
To examine target-centric binding characteristics of de novo generated molecules, we conducted a molecular docking analysis (please see “Methods” for details) on the de novo molecules generated by DrugGEN and competing targeted molecule generation methods (i.e., RELATION, TRIOMPHE-BOA, and ResGen). We also employed the real AKT1 inhibitors as the positive control, and baselines (i.e., DrugGEN-NoTarget, MLP, MLP-NoTarget, and randomly selected ChEMBL molecules) as the negative control. We utilised the AKT1 crystal structure (PDB id of the model: “4GV1”)[42] as our template. Figure 2A displays bar plots of median docking scores (i.e., binding free energies - $\delta$G in kcal/mol) of the top 10% of molecules with the best binding properties from each group/model, where a lower binding free energy value indicates a stronger binding affinity. Predicted binding free energies for all molecules are given as histograms in Figure S3. The median docking scores of DrugGEN and positive/negative controls were already provided as part of the ablation study, above (Table 2).
In Figure 2A, percentages shown on top of each bar represent docking performance of generative models in comparison to the real AKT1 inhibitors. For this, we applied min-max normalisation to the median docking score of each model (parameters: maximum=-9.41 and minimum=-3.75 kcal/mol, which are the median of the top and bottom 10% docking scores of real AKT1 inhibitors, respectively). DrugGEN achieved the highest performance with 81.64% and -8.37 kcal/mol (Figure 2A), exhibiting a noteworthy contrast to other target-based models, especially TRIOMPHE-BOA, and ResGen, as well as the non-targeted model, DrugGEN-NoTarget. Furthermore, the difference between the DrugGEN and RELATION models (i.e., -8.37 and -8.10 kcal/mol, respectively) was found to be statistically significant according to the Mann-Whitney U rank test (p-value < 0.0001). Additionally, we measured the docking binding free energy of the capivasertib-AKT1 complex (i.e., the bound ligand in our template PDB structure: 4GV1) at -9.52 kcal/mol, as well as other experimentally proven inhibitors that reached clinical trials, namely ipatasertib and uprosertib, at -9.68 and -8.57 kcal/mol, respectively. The DrugGEN model was able to generate 38 molecules that surpass the binding score of the native ligand capivasertib, which are shown in Figure S4.
As an overall evaluation, DrugGEN was successful in its target-specific molecular design routine. After this stage of the study, we concentrated on further investigating DrugGEN’s de novo molecules and selecting a subset to propose as new potential inhibitors for AKT1.
@@ 2.6 Exploration of Molecules via Dimensionality Reduction
We visualized both real and de novo molecules from the previous analyses using UMAP [50] and t-SNE [51] algorithms. Figure 2B and C present the UMAP (parameters: n-neighbors=50, min-dist=0.8 and metric=”dice”) and t-SNE (parameters: perplexity: 500, iteration: 2000) projections, respectively, of a randomly selected subset of de novo molecules from DrugGEN and DrugGEN-NoTarget models (1,000 from each), real AKT1 inhibitors (all molecules used in model training); and 50,000 ChEMBL molecules from the training dataset. In these plots, each dot represents a molecule, colours indicate their source, and the Euclidean distances roughly indicate structural similarities based on Tanimoto applied on descriptor-based molecular fingerprints. The molecules are represented by the MACCS (Molecular ACCess System) descriptors, which is a 166-D set of binary fingerprints concerning the presence or absence of predefined structural features [52].
In the UMAP and t-SNE plots in Figure 2B and C, molecules generated by the DrugGEN model are clustered together with real AKT1 inhibitors, indicating a higher structural resemblance (in terms of topological fingerprint-based similarities). On the other hand, molecules generated by DrugGEN-NoTarget exhibit a nearly homogeneous distribution across the entire plane, due to the non-targeted generation approach. Our conclusion here is that the DrugGEN model has learned the structural distribution of the AKT1 inhibitor molecules.
2.7 Analysing Molecules with Drug-likeness Filters
To assess their drug-related properties, de novo molecules generated by DrugGEN are subjected to the following filtering steps: (i) Lipinski’s Rule of 5 [53], (ii) Veber’s rules [54], and (iii) the pan-assay INterference compoundS (PAINS) [55] filter,which identifies and eliminates false positive molecules in biological screening assays such as redox cyclers, toxoflavins, polyhydroxylated natural phytochemicals, etc. In this analysis, we utilised the same 10,000 de novo generated molecules subjected to the benchmarking analysis. In the end, 41.15% (4,115 out of 10,000) of the de novo molecules successfully passed through the filters. Subsequently, we subjected ChEMBL and AKT1 molecules to the same filters for comparison. It is observed that 57.18% (1,342 out of 2,347) of AKT1 inhibitors and 50.80% (50,802 out of 100,000) of randomly sampled ChEMBL dataset molecules have passed, providing roughly similar percentages to DrugGEN molecules.
@@ 2.8 Deep Learning-based Bioactivity Prediction
To evaluate the target interaction-related properties of de novo molecules in parallel to molecular docking, we carried out a deep learning-based drug/compound–target interaction (DTI) prediction analysis with the same de novo molecule dataset (against the AKT1 protein), using our previously developed DEEPScreen system [56]. DEEPScreen is a supervised discriminative deep learning system composed of target-centric bioactivity prediction models that classify input molecules as either active (interacting) or inactive (noninteracting) against the selected protein (Figure 2D) (please see “Methods” for details). We preferred DEEPScreen for this analysis since (i) it is straightforward to prepare input data and train/test a target-specific DTI prediction model, (ii) it only requires readily available data, (iii) it has high predictive performance as displayed in its respective article [56]. DEEPScreen is completely separate/ independent from the DrugGEN system regarding the modelling approach, datasets and output. Therefore, it is ensured that the DTI prediction analysis is objective. Out of 10,000 molecules obtained from the DrugGEN model, DEEPScreen predicted 5,700 of them as active, and 373 of these active predictions received an extraordinary confidence score of $\geq$ 0.85 (the complete confidence score histogram of the active predictions is given in Figure S5). We also analysed 2,347 real AKT1 inhibitors, 10,000 random ChEMBL compounds and 10,000 DrugGEN-NoTarget molecules with DEEPScreen for comparison, which received 777, 168 and 138 high confidence ($\geq$ 0.85) active predictions against the target protein, respectively. The high ratio of real AKT1 inhibitors can be deceiving here since most of these “active” predicted molecules were already presented in the positive training or validation datasets of the DEEPScreen model. As a result, the model was highly biased towards predicting these molecules as active. Considering the two non-targeted molecule groups (i.e., DrugGEN-NoTarget and random ChEMBL molecules), the difference in the number of high-confidence active predictions (373 vs. 168 and 138) indicates the effectiveness of DrugGEN. In accordance with the results of the docking analysis, these findings provide additional information on DrugGEN’s ability to generate target-based molecules.
Figure 2: Exploration of de novo molecules via downstream analysis. (A) Bar plots displaying the median of binding free energies measured in the docking analysis of de novo molecules generated by DrugGEN, DrugGEN-NoTarget, RELATION, TRIOMPHE-BOA, and ResGen models (molecules are docked into the binding pocket of the AKT1 protein structure). The whiskers on the bars represent the standard error of the median (median values are utilised due to the non-normal distribution of data). The percentages given above each bar represent docking scores normalised with respect to the scores of real AKT1 inhibitors, with the top real AKT1 inhibitors set at 100%. (B-C) 2-D visualisation of the molecules generated by DrugGEN and DrugGEN-NoTarget models, real AKT1 inhibitors and randomly selected ChEMBL molecules via UMAP (B) and t-SNE (C) projections. (D) Deep learning-based bioactivity prediction analysis. Top: the work-flow of DEEPScreen, which uses 2-D pixel-based (image) structural representations of molecules as input and processes them via deep convolutional neural networks [56]. Middle: the DEEPScreen AKT1 model is trained with binarised experimental bioactivity data points of the AKT1 protein (obtained from ChEMBL binding assays) and subjected to DrugGEN’s de novo molecules in the inference/prediction mode. Right: bioactivity prediction results, i.e., the number of high-confidence active predictions among the real AKT1 inhibitors, randomly selected ChEMBL molecules, and DrugGEN and DrugGEN-NoTarget molecules.
2.9 Final Candidate Molecules Selection
@@
We manually selected the 30 most promising de novo molecules (from molecules with docking scores < -8 kcal/mol and DEEPScreen predictions as “active”) via expert curation and presented them as our best candidates to target AKT1 (Figure 3). Figure S6 shows a subset of those 30 de novo molecules alongside the real AKT1 inhibitory molecules (from the training dataset) most similar to them (two real molecules for each de novo design). As observed from Figure S6, de novo molecules are structurally diverse even from their most similar counterpart in the training set, yet these molecules obtained convincing docking scores (i.e., < -8 kcal/mol) against AKT1.
We showcase one molecule among the 25 (shown with the identifier: “$MOL\_01\_027820$” in Figure 5), which can be denoted as a “Pyrrolo[1,2-a]pyrimidin-4(1H)-one” derivative. Figure 4A displays the reference crystal complex structure of AKT1 with the ligand capivasertib (model PDB id: “4GV1”) [42], for which the binding free energy was measured as -9.517 kcal/mol in the molecular docking analysis (on the same ligand conformer). Figure 4B shows the best docking pose of $MOL\_01\_027820$ with the binding free energy -9.686 kcal/mol obtained via the same docking protocol. This molecule is investigated further below to comprehend its binding-related dynamic properties.
@@
Figure 3: 30 promising de novo molecules to effectively target AKT1 protein (generated by DrugGEN model), selected via expert curation from the dataset of molecules with sufficiently low binding free energies (< -8 kcal/mol) in the molecular docking experiment and deep learning-based DTI predictions (by DEEPScreen [56]) as “active”.
2.10 Investigation of Binding Properties of the de novo Showcase Molecule
The reference co-crystal complex structure “AKT1 - capivasertib” (Figure 4A) and the complex model obtained from the docking of the selected de novo molecule “AKT1 - $MOL\_01\_027820$” (Figure 4B) share the binding pattern with Ala177, Lys179 and Thr291. Hydrophobic interactions are more prominent in the AKT1 - $MOL\_01\_027820$ complex, whereas in the crystal structure hydrogen bonds are evident. This may contribute to the difference in binding affinities. The replacement of the adenine ring of the ATP with various heterocyclic moieties resulted in selective and potent inhibition of kinases and this mode of binding constitutes more than 90% of kinase inhibitors [57]. These inhibitors bind the same binding pocket as ATP and mimic adenine group’s interactions with the hinge residues. Here we see a similar orientation with capivasertib and $MOL\_01\_027820$ in the AKT1 kinase binding site. However, the hydrogen bond interactions with the backbones of the hinge residues were not directly observed in the docking-derived binding pose of $MOL\_01\_027820$ (Figure 4B). To further evaluate the relaxation and adaptation of the binding site and to reveal binding characteristics of capivasertib and $MOL\_01\_027820$, 500 ns long molecular dynamics (MD) simulations were conducted with four copies (please see “Methods” for details).
The output of MD simulations are shown in Figure 4 panels C to I. The AKT1 structure mostly consists of $\alpha$ helices, and visual inspection of the simulations displayed the conservation of secondary structures during the simulations of both complexes. As can be seen in RMSD plots (Figure 4 panels F to I), both simulations (together with their simulation copies) reached stable binding patterns. Hydrogen bonding networks between the hinge domain residues Glu228 and Ala230 occurred with high ratios in all simulations. Van der Waals interactions were also observed with the hydrophobic moieties of the surrounding residues. The gatekeeper residue Met227 remained closed and supported the residence of the ligand within the binding cavity for both complexes. Additionally, Lys179 in $\beta$-sheet continued to form a salt bridge with Glu234 in the $\alpha$C-helix located at the N terminal domain of AKT1. This was a supporting factor to keep the ligand within the binding site. With the help of this ionic interaction, the protein stayed in DFG-in and $\alpha$C-helix in conformation in the simulations.
Capivasertib remained stable during all simulations (RMSD <1, Figure 4 panels F and H). The interaction analysis of capivasertib showed that hydrogen bonds with the hinge domain (via Glu228 -99%- and Ala230 -99%- residues) were conserved during the simulation. An additional hydrogen bond was seen with Glu278 (74%). $MOL\_01\_027820$ showed two different binding patterns (pattern-1 and -2, shown in Figure 4 panels D and E, respectively). Those patterns were seen in two individual copies of the simulations. Pattern-1 was observed in the second and the third replica simulations and pattern-2 was observed in the first and the fourth replicas (Figure 4G). The first pattern was similar to the docking-derived starting conformation ($RMSD_{average}$ = 1.2, Figure 4 panels D, G, and I) and relatively close to the binding pattern with the compound capivasertib. However, the second pattern showed 110.2 degrees of dihedral tilting ($RMSD_{average}$ = 3.1, Figure 4 panels E, and G) following atom number 15 and above (Figure 4I) but still conserved the essential interactions with the hinge domain and remained within the binding cleft and conserved the protein’s general conformation. Further comments can be found in Supplementary Material (S5).

### v7

@@ 2.4 Ablation study
In an extensive ablation study, we compared several variations of our model to assess the effectiveness of the graph transformer architecture and the modified attention mechanism in DrugGEN for target-specific molecule generation. Specifically, we evaluated the default DrugGEN model and its non-targeted variant (DrugGEN-NoTarget) against baseline models that use simpler architectures—multi-layer perceptron (MLP) or graph convolutional network (GCN) modules—in both the generator and discriminator modules. Additionally, we analysed the attention mechanism to investigate various strategies for incorporating edge information (i.e., atomic bonds). For each model variant, 10,000 de novo molecules were generated and evaluated using the same fundamental generative metrics and drug-likeness scores as in Tables 1 and 2. Overall, the ablation study confirms that the use of graph transformer blocks with the modified attention, involving an additional multiplication by a shifted adjacency matrix, i.e., $Att\bigodot A_{m}\bigodot(A_{m}+1)$, significantly enhances the model’s ability to generate valid, diverse, and novel molecules with high drug likeness, demonstrating superior performance compared to the MLP- and GCN-based baselines and attention variants. Please see Supplementary Material (S6) and Table S1 for detailed descriptions and complete results.
@@ 2.5 Exploration of Molecules via Dimensionality Reduction
We visualized both real and de novo molecules from the previous analyses using UMAP [58] and t-SNE [59] algorithms. algorithms (please see Methods for details). In the UMAP and t-SNE plots in Figure 2B, molecules generated by the DrugGEN model are clustered together with real AKT1 inhibitors, indicating a higher structural resemblance (regarding topological fingerprint-based similarities). On the other hand, molecules generated by DrugGEN-NoTarget exhibit a nearly homogeneous distribution across the entire plane due to the non-targeted generation approach. Our conclusion is that the DrugGEN model has learned the structural distribution of the AKT1 inhibitor molecules.
@@ 2.6 Deep Learning-based Bioactivity Prediction
To evaluate the target interaction-related properties of DrugGEN’s de novo molecules in parallel to molecular docking, we carried out a deep learning-based drug/compound–target interaction (DTI) prediction analysis against the AKT1 protein, using our previously developed DEEPScreen system [57]. DEEPScreen is a supervised discriminative deep learning system composed of target-centric bioactivity prediction models that classify input molecules as either active (interacting) or inactive (non-interacting) against the selected protein (Figure 2C). For this analysis, we trained and tested a DEEPScreen model using the experimentally proven active and inactive small molecules of AKT1 (please see “Methods” for details).
When tested on the scaffold-split hold-out dataset composed of 484 known AKT1 binder molecules, DEEPScreen AKT1 model accurately predicted 475 as active molecules based on a standard confidence score threshold of 0.5 (i.e., at least half of the augmented molecule images are predicted as active), resulting in precision: 0.96, recall: 0.94, F1-score: 0.95, and MCC: 0.91, which was considered highly satisfactory. Among 484 known AKT1 binders, 421 were classified with a high confidence score ( 0.85), indicating the reliability of the model.
We tested 10,000 de novo molecules generated by the DrugGEN model using DEEPScreen, which resulted in 1,382 active predictions based on the standard confidence threshold ($\geq$0.5). Of these, 748 molecules were predicted with high confidence ($\geq$ 0.85). A complete histogram of the confidence scores for active predictions is provided in Figure S6. DrugGEN molecules, on average, exhibit a lower average molecular similarity (Tanimoto) to the training set of DEEPScreen when compared with the hold-out test dataset (0.309 and 0.637, respectively), which may explain why the proportion of de novo molecules predicted to be positive is not as high. We also analysed 10,000 random ChEMBL compounds and 10,000 DrugGEN-NoTarget molecules with DEEPScreen for comparison, which received 122 and 69 high-confidence ($\geq$ 0.85) active predictions against AKT1, respectively. Considering the two non-targeted molecule groups (i.e., DrugGEN-NoTarget and random ChEMBL molecules), the difference in the number of high-confidence active predictions (748 vs. 122 and 69) indicates the effectiveness of DrugGEN in generating target-based molecules.
2.7 Selection of thirty promising molecules
We manually picked 30 promising de novo molecules (mainly from molecules with docking scores < -8 kcal/mol and DEEPScreen predictions as “active”) via expert curation and presented them as our best candidates to target AKT1 (Figure 3). Many of these selected molecules can be considered readily synthesisable and exhibit favorable drug-like properties as indicated by their SA scores (< 6 [60]), and QED scores (> 0.5 [61]), while the others display less conventional features, such as increased polarity or molecular weight. Compliance with Lipinski’s Rule of Five was also confirmed, with hydrogen-bond donors ($\leq$ 5) and acceptors ($\geq$ 10) within acceptable limits, further supporting their potential as drug candidates.
@@ 2.7 Selection of thirty promising molecules
Figure S7 shows a subset of those 30 de novo molecules alongside the real AKT1 inhibitory molecules (from the training dataset) most similar to them (two real molecules for each de novo design). As observed from Figure S7, de novo molecules are structurally diverse even from their most similar counterpart in the training set. Yet, most of these molecules obtained convincing docking scores (i.e., < -8 kcal/mol) against AKT1.
@@
Figure 3: 30 promising de novo molecules to effectively target AKT1 protein (generated by DrugGEN model), selected via expert curation from the dataset of molecules with sufficiently low binding free energies (< -8 kcal/mol) in the molecular docking experiment and deep learning-based DTI predictions [57]) as “active”.
2.8 Experimental validation of de novo molecules for AKT1 inhibition
Following the computational analyses, chemical synthesis and in vitro enzymatic assays were conducted for experimental validation. Among the 30 promising de novo compounds given in Figure 3, five structurally diverse compounds were selected to be synthesised (Figure 4A). These compounds were chosen based on their economic synthetic accessibility and low synthesis and purification turnaround times. Figure 4B shows the best docking poses of Molecules 1 ($MOL\_02\_045762$), 2 ($MOL\_02\_045795$), 3 ($MOL\_02\_045758$), 4 ($MOL\_02\_045609$), and 5 ($MOL\_02\_045627$) with binding free energies of -8.948, -7.686, -8.655, -8.383, and -5.341 kcal/mol, respectively. Additionally, Figures 4A and 4B include the reference crystal structure ligand, capivasertib, for comparison, with a docking score of -9.517 kcal/mol. Details about the docking results of synthesised molecules are provided in Supplementary Material (S9).
Selected compounds were synthesised by commercial organic chemistry laboratories. The analytical proof for the purities of the compounds can be found in Figures S12-16. Subsequently, in vitro enzymatic assays were conducted on these compounds, which rely on transferring a radioactive 33P-labeled phosphate group from ATP to the target kinase (AKT1) substrate. The study yielded two out of five chemically different compounds, showing IC50 values of 1.89 µM (Molecule 1) and 48.6 µM (Molecule 2) against AKT1 (Table S3). However, Molecules 3, 4, and 5 exhibited insufficient activity up to 100 µM. Corresponding dose–response curves are displayed in Figure S17. To rationalise the experimental results, molecular dynamics simulations were performed and explained below.